OpenStudy (anonymous):

A truck with a mass of 3000kg is accelerated uniformly from rest to 6m/s1 over a distance of 600m up an incline of 1:30. The resistance against motion is 200N. Calculate the component of gravitational force parallel to the plane. Please show your calculations.

OpenStudy (anonymous):

if you had the angle of the incline, cant you use 29400 x sin.angle?

OpenStudy (anonymous):

3000 x 9.8 = 29400N

OpenStudy (ujjwal):

 The gravitational force parallel to plane will always be equal to mg\(\sin \theta\). where \(\theta\) is the angle of inclination.